OpenStudy (anonymous):

how many atoms are in 165 g of calcium

OpenStudy (anonymous):

\[n=\frac{ m }{ M }=\frac{ N }{ N _{A} }\]

OpenStudy (anonymous):

you can use any two sets of the equation depending on what you are given..... n=no. of moles m=mass M=Molar mass (mass of 1 mole of the atom, which can be found on the periodic table) N=No. of particles (atoms, molecules, ions,....) Na=Avogadro's constant (6.02 x 10^23)

OpenStudy (anonymous):

\[\frac{ 165 }{ 40 }=\frac{ N }{ 6.02 \times 10^{23} }\] \[N=2.48325 \times 10^{24}\]